### v1

@@ 3.2.2 Transient state
Figure 7: The mechanism of producing AP (using numbers referring to the case of squid shown in Table 1). In the resting state, the inside thermal potentials follow the green paths. When the rush-in of $Na+$ ions suddenly increases the internal concentration to $[200]$, the potential on the internal side of the membrane increases to $+90\ mV$, and the neuron must decrease its $K^{+}$ concentration to restore the resting potential, or release an action potential through the AIS.
In a special type of invasion, when suddenly a large amount of $Na^{+}$ rushes in (i.e., the trigger is external, but the ions derive from a biological process), the current throughput of the "resting ion channels" is not sufficient anymore: the internal $Na^{+}$ concentration drastically increases, and so does the membrane’s potential. At that point, the neuron is out of balance: although most of the $Na^{+}$ ions can flow out through the AIS, the system also changes its $K^{+}$ concentration through the limited capacity "resting ion channels" until the balanced state regained. This latter effect must not be mismatched with the capacitive current, which (due to its changed direction) causes the phenomenon known as "hyperpolarization".
Compare the figure to Fig. 7.2 in [1]. Notice the two-sided negative feedback effect: a change in the value of the "command potential" or "command concentration" changes the other value. In the case of voltage invasion, the ratio of the ion concentrations also changes. In this case, the feedback amplifier is represented by the delicate balance of the electric and thermodynamic force fields. In the case of clamping, another (electric) feedback amplifier is introduced into the system and they work against each other. The neuron attempts to restore its balance without external potential (it cannot distinguish the foreign potential from its membrane potential), and changes its concentrations accordingly. A vital difference is the speed of the feedback amplifier. The electrical one works with electric speed, while the thermodynamic one with the several orders lower thermodynamic speed. When switching (or changing) the external ("command") voltage, the neuron must first get into its "steady state". It takes time. "The command potential, which is selected by the experimenter and can be of any desired amplitude and waveform" [1]. However, if the measured data readings are not from a steady state, the measurement is wrong: the different amplifier speeds (different current speeds) surely distort the measured value.
@@
On the left side, two positive ions are in the solution, and the membrane potential (due to charge separation) rises toward the outside segment; the ions cannot go "uphill". Similarly, the $Cl^{-}$ ions cannot go uphill (due to their negative charge, the same potential is also rising for them). This single membrane potential keeps all ions in their segments. However, the $Ca^{2+}$ ions can travel "downhill". Calcium pumps needed: although in minimal concentration, $Ca^{2+}$ ions flow into the intracellular segment, and they must be pumped out. (BTW: this is likely also the reason why measuring $Ca^{2+}$ concentration is inaccurate.)
@@
As we discussed above, the concentrations and voltages in the extracellular segment, with its vast amount of ions, remain unchanged in resting and transient states. Follow the red path: when $Na^{+}$ ions rush into the intracellular segment; they increase the concentration from $[50\ mM]$ to (say) $[200\ mM]$ (from point (Rest) to point (AP)). This sudden change increases the thermal contribution of $Na^{+}$ potential to $[-20\ mV]$, but because the total charge on the membrane significantly increases, the resulting voltage on the intracellular side of the membrane increases to $[90\ mV]$. The neuron could decrease its potential to the resting value if it could decrease $K^{+}$ concentration to $[100\ mM]$. However, the ion delivering capacity of the "resting ion channels" is not sufficient: they are sized to maintain the resting state. (Furthermore, the electric contribution appears instantly, while the thermodynamic one, due to the finite speed of ions, only with some delay [7].) Follow the dashed orange path: the large amount of rush-in $Na^{+}$ ions drastically increases the surface concentration on the membrane, so the potential suddenly changes to $[90\ mV]$. Follow the dotted orange path: the electric field across the membrane suddenly changes its direction from positive to negative (see the slopes of the solid and dotted orange lines) so that the excess positive ions can move "downhill" out of the intracellular space through the high-capacity ion channel array (AIS) at the end of the neuron. Its internal end is at the high membrane potential, while the external end is at the low potential of the extracellular segment.
@@
The set-point on the extracellular side remains fixed. On the intracellular side, at the beginning of the AP, the actual voltage is above that of the set-point on the extracellular side, so at the two ends of the AIS, a positive driving force moves the ions toward the downstream neuron. Given that positive $Na^{+}$ ions leave the intracellular segment, its potential continuously decreases. At some point, the voltage drops below the set-point of the extracellular side and the electric field across the membrane reverses. The driving force also decreases, and a low-intensity current restores the balanced state, to the previous set-point. When measuring at the AIS, one observes that the potential rises suddenly, then decreases below the set-point (hyperpolarizes the membrane), then a decreasing current (flowing in the opposite direction due to the changed slope of the resulting potential (the changed electric field across the membrane), but comprising the originally rushed-in $Na^{+}$ ions) slowly restores the resting potential. Here comes to light the finite speed of ions again. The ions can follow the potential changes with a delay due to their finite speed, causing an apparent delay in the time course of the AP’s current. Until they travel, the field at the AIS changes, but the finite speed of the ions cannot deliver that field in time. The current "overshoots" then relaxes. This is the normal operation of an $RC$ circuit; the capacitive current is misinterpreted as delayed $K^{+}$ current.
3.3 Mathematics of the transient process
During regular operation, a very intense current pulse is created: "A small flow of ions carries sufficient charge to cause a large change in the membrane potential. The ions that give rise to the membrane potential lie in a thin ($<1\ nm$) surface layer close to the membrane" [26] (see the dotted line in Fig. 4; notice that the horizontal scale on the figure spans a few $nm$). For this, a considerable number of ion channels are used. When the trigger signal is received, the gated ion channels in the membrane’s wall open quickly and only for a brief period. The rushed-in ions create a large offset potential in the layer proximate to the membrane’s surface (from an electrical point of view, a condenser) that drives an intense current through the always-open ion channel array AIS (from an electrical point of view, a resistor) at the beginning of the axon. The slow current on the membrane’s surface mimics a condenser: in the picture of classical electricity, the charges stay inside the distributed condenser while they travel on its surface. In classical electronics, the condenser is sizeless.
Those two elements form a serial $RC$ circuit that acts as a damped oscillator and produces a single current pulse, known as AP. Neuroscience mistakenly claims it is a parallel $RC$ circuit. As discussed in the previous subsection, there are ’resting’ ion channels in the membrane’s wall, and they are in parallel with the resistor. However, their number and intensity are insignificant for producing AP. In this action, the gated ion channels in the membrane’s wall produce an instant pulse that appears as a peaked current time course due to the slowness of the charge carriers. The output current flows as long as the offset potential persists.
@@ 3.3 Mathematics of the transient process
The process can be modeled as follows: a sudden voltage gradient appears on the membrane (as a discrete capacitor), and a current flows out through the serially connected ion channel array (as a discrete resistor). The neuron forms a simple serially (not parallelly, as assumed by Hodgkin and Huxley [28] and mistakenly claimed by neurophysiology) connected $RC$ oscillator. For the physical and mathematical description, see [30], section 4; furthermore, for further details [14]. As we discussed there, the capacitive current, which per definitionem changes its direction and so generates an opposite voltage on the resistor represented by the AIS, perfectly describes the so-called "hyperpolarization". It is not the effect of a $K^{+}$ current: the resting ion channels do not have enough transmission capacity, although some $K^{+}$ current surely exists.

### v2

@@ 3.1.2 Transient state
Figure 10: The mechanism of producing AP (using numbers referring to the case of squid shown in Table 1). In the resting state, the inside thermal potentials follow the green paths. When the rush-in of $Na^{+}$ ions suddenly increases the internal concentration to $[200]$, the potential on the internal side of the membrane increases to +90mV+90\ mV. The neuron must decrease its concentrations to restore the resting potential, mainly by releasing an action potential through the AIS.
In a special type of invasion, when suddenly a large amount of $Na^{+}$ rushes in (i.e., the trigger is external, but the ions derive from a biological process), the current throughput of the "resting ion channels" is not sufficient anymore: the internal $Na^{+}$ concentration in the proximal layer drastically increases, and so does the membrane’s potential. At that point, the neuron is out of balance: although most of the $Na^{+}$ ions can flow out through the AIS, the system also changes its $K^{+}$ concentration through the limited capacity "resting ion channels" until the balanced state is regained. This latter effect must not be mismatched with the capacitive current, which (due to its changed direction) causes the phenomenon known as "hyperpolarization".
Compare the figure to Fig. 7.2 in [1]. Notice the two-sided negative feedback effect: a change in the value of the "command potential" or "command concentration" changes the other value. In the case of voltage invasion, the ratio of the ion concentrations also changes. In this case, the feedback amplifier is represented by the delicate balance of the electrical and thermodynamic force fields. In the case of clamping, another (electrical) feedback amplifier is introduced into the system, and they work against each other. The neuron attempts to restore its balance without external potential (it cannot distinguish the foreign potential from its membrane potential), and changes its concentrations accordingly. A vital difference is the speed of the feedback amplifier. The electrical one works with electrical speed, while the thermodynamic one with several orders lower thermodynamic speed. When switching (or changing) the external ("command") voltage, the neuron must first get into its "steady state". It takes time. "The command potential, which is selected by the experimenter and can be of any desired amplitude and waveform" [1]. However, if the measured data readings are not from a steady state, the measurement is wrong: the different amplifier speeds (different current speeds) surely distort the measured value.
@@
On the left side, two positive ions are in the solution, and the membrane potential (due to charge separation) rises toward the outside segment; the ions cannot go "uphill". Similarly, the $Cl^{-}$ ions cannot go uphill (due to their negative charge, the same potential is also rising for them). This single membrane potential keeps all ions in their segments. However, the $Ca^{2+}$ ions can travel "downhill". Calcium pumps needed: although in minimal concentration, $Ca^{2+}$ ions flow into the intracellular segment, and they must be pumped out (BTW: this is likely also the reason why measuring $Ca^{2+}$ concentration is inaccurate).
@@
As we discussed above, the concentrations and voltages in the extracellular segment, with its vast amount of ions, remain unchanged in resting and transient states. Follow the red path: when $Na^{+}$ ions rush into the intracellular segment. They increase the concentration from [50mM][50\ mM] to (say) [200mM][200\ mM] (from point (Rest) to point (AP)). This sudden change increases the thermal contribution of $Na^{+}$ potential to [20mV][-20\ mV], but because the total charge on the membrane significantly increases, the resulting voltage on the intracellular side of the membrane increases to [90mV][90\ mV]. The neuron could decrease its potential to the resting value if it could decrease the $K^{+}$ concentration to [100mM][100\ mM]. However, the ion-delivering capacity of the "resting ion channels" is not sufficient: they are sized to maintain the resting state. (Furthermore, the electrical contribution appears instantly, while the thermodynamic one, due to the finite speed of ions, only with some delay [8].) Follow the dashed orange path: the large amount of rush-in $Na^{+}$ ions drastically increases the surface concentration on the membrane, so the potential suddenly changes to [90mV][90\ mV]. Follow the dotted orange path: the electric field across the membrane suddenly changes its direction from positive to negative (see the slopes of the solid and dotted orange lines) so that the excess positive ions can move "downhill" out of the intracellular space through the high-conductance ion channel array (AIS) at the end of the neuron. Its internal end is at the high membrane potential, while the external end is at the low potential of the extracellular segment.
@@
The setpoint on the extracellular side remains fixed. On the intracellular side, at the beginning of the AP, the actual voltage is above that of the setpoint on the extracellular side, so at the two ends of the AIS, a positive driving force moves the ions toward the downstream neuron. Given that positive $Na^{+}$ ions leave the intracellular segment, its potential continuously decreases. At some point, the voltage drops below the setpoint of the extracellular side, and the electric field across the membrane reverses. The driving force also decreases, and a low-intensity current restores the balanced state to the previous setpoint. When measuring at the AIS, one observes that the potential rises suddenly, then decreases below the set-point (hyperpolarizes the membrane), then a decreasing current (flowing in the opposite direction due to the changed slope of the resulting potential (the changed electrical field across the membrane), but comprising the originally rushed-in $Na^{+}$ ions) slowly restores the resting potential. Here comes the finite speed of ions again. The ions can follow the potential changes with a delay due to their finite speed, causing an apparent delay in the time course of the AP’s current.
Models in neuroscience (as reviewed in [66]) almost entirely ignore these aspects. In our physical model, we see that the measurable membrane potential and current change in the function of the ions’ speed, the concentration, and its time derivative; furthermore, all mentioned quantities depend on the effective potential. As Eq.(3) shows, the "thermodynamic electric field" increases as the temperature increases, and raises the setpoint. As Eq.(27) shows, the current decreases as the temperature increases, in line with the experimental evidence [50, 67]. The amplitude of the action potential is decreased, given that the setpoint has increased and its duration is reduced, given that the current decreases.
4 The Action Potential
As discussed, a smaller number of ion channels is distributed over the surface of the membrane, and the overwhelming majority of ion channels is concentrated in the AIS.
4.1 The physical process
In the resting state, we have a condenser that has some low-conductance resistors (the resting ion channels) and a very low intensity (nearly constant) distributed current flows, under the effect of a slightly varying (nearly constant) membrane voltage. Due to the membrane potential, a current through the always-open channels flows on the dendritic surface and the resting channels’ conductivity represents a drain with sufficient transmission. The resting current practically does not reach the AIS: the ion channels along the current’s path practically "shunt" the current. This is a static state that classic neuroscience assumes: some static current in and some static (leakage) current out; no significant gradient. In this state, the correct model is a condenser with a parallelly switched resistor; an integrator-type $RC$ circuit, see Table 2 (although when approaching the threshold voltage, the AIS plays some role. The smaller gradient changes, such as subthreshold excitations, produce "mini-AP waves" [15], providing direct experimental proof that in this state the parallel $RC$ circuit is not correct). Some perturbations must be counterbalanced, but their amount does not exceed a predefined threshold.
In the transient (perturbed) state, the process variable (the membrane’s potential) exceeds the threshold. That event acts as a fast trigger signal. The amount of charged particles (and so: the membrane’s voltage) suddenly increases, and the created gradient drives a current. When $Na^{+}$ ions rush into the intracellular layer, they roughly increase the overall concentration and the potential in that thin layer. All other ions, including $K^{+}$, also feel a driving force. The targeted membrane potential is set electrically, and the driving gradients may behave controversially during the transient period. The actual voltage gradient may temporarily direct the chemical gradients in opposite directions. Phenomena such as ’over-relaxation’, in this sense, similar to ’hyperpolarization’, may happen.
Our results align with the observation (see caption of 11.22 in [45]), that the significant processes occur in a thin layer of the electrolyte proximal to the membrane surface. The amount of unbalanced ions is in the range of $10^{7}$, and so is the amount of rush-in ions. In addition, those ions on the high-concentration side rush into the low-concentration side and cause a significant change in the membrane potential. Their absolute amount is small compared to the total number of ions in the cell, but it is significant compared to the number of unbalanced ions in that layer. However, the layer itself can also be modeled as having just a few ions under their mutual repulsion on the surface or in a few atomic layers on top of each other, depending on the concentration.
Suppose we assume that at the beginning of an AP a large amount of $Na^{+}$ ions are transferred from the extracellular to the intracellular side. In that case, their concentrations change from $145/15$ to $45/115$ and the corresponding thermodynamic voltage contribution changes from +61[mV]+61\ [mV] to 25[mV]-25\ [mV] (i.e., altogether a 86[mV]86\ [mV] sudden increase in membrane potential). For the $Na^{+}$ ions, the resultant potential changes from +4[mV]+4\ [mV] to 83[mV]-83\ [mV]. That means when the AP begins, the "Na-K pump" stops (if the resulting potential provides the driving force for the exchange pump. The only way for the neuron to remove the excess $Na^{+}$ ions is to generate a current toward the AIS (where the other end of the ion channels remained at the extracellular potential), until the $Na^{+}$-specific driving force disappears. The $K^{+}$-specific driving force, changes from 90[mV]-90\ [mV] to 176[mV]-176\ [mV], i.e., the $K^{+}$ intake gets more intensive, misleading researchers into believing that it causes the observed hyperpolarization. However, as Fig. 3 in [15] demonstrates, it occurs instantly, not with a delay; furthermore, the "leakage current" and the stimulus current are negligible. The low extracellular $K^{+}$ concentration plus the low number of channels in the membrane’s wall do not enable a significant increase in the intracellular $K^{+}$. (Given that complex changes occur, including changes in the electric potential that also change the concentrations, different waves start; the statement is not strictly valid.)
As the current flows, the thermodynamic force decreases, and at some point, the resultant force changes its sign; so does the direction of the current. This is seen as the "condenser current" (and mis-identified as an intense $K^{+}$ current). The ions move slowly, and due to the enormous electrostatic repulsion, their "electrostatic fluid" must be continuous. A special "damped wave" is formed (with all secondary mechanical, optical, etc. effects; see section 4.4) and the condenser current reverses. Electricity interprets the phenomenon as "capacitive current" in the equivalent circuit, and thermodynamics interprets it as changed chemical concentration (or the negative amplitude of the damped elastic oscillation). Our discussion adds that the electric repulsion generates a pressure with a shock wave in the neuron’s closed volume.
Figure 11: A summary of producing neuronal AP as a 3-D figure. The thick solid 3-D line shows how the color-coded voltage gradient drives the membrane potential in the function of the time. The three dashed diagram lines are the projections of the spatial curve to the boundary planes of the cube. These 2-D plots are delivered by the physiology; the figure explains how those parameters are interrelated. The red diagram line shows the voltage gradient produced by the rush-in of the $Na^{+}$ ions, see also the middle inset in Fig. 12. It produces the blue diagram line AP, see also the bottom inset in Fig. 12. The interdependence of the gradient and the voltage it produces leads to Fig. 14, left side, that allows to calculate the thermodynamical characteristics, right side, of neuronal operation.
The neuronal AP generation can be entirely described by the 3-D line in Fig. 11. The three dashed diagram lines are drawn in three perpendicular planes; see the figure’s caption and the discussion below for the details.
4.2 The oscillator model
Despite the early warning that ’it was not possible to separate the change into resistance and capacity components[31], a commonly accepted truism was that neurons, in some sense, behave as electric oscillators. HH introduced the idea explicitly that the electrically equivalent circuit of a neuron is an $RC$ oscillator. We must also add the invention from about three decades later: "Neurons ensure the directional propagation of signals throughout the nervous system. The functional asymmetry of neurons is supported by cellular compartmentation: the cell body and dendrites (somatodendritic compartment) receive synaptic inputs, and the axon propagates the action potentials that trigger synaptic release toward target cells. Between the cell body and the axon sits a unique compartment called the axon initial segment. The AIS was first described 50 years ago [i.e., nearly two decades after HH published their study], and its molecular composition and organization have been progressively elucidated during the following decades. …. Recent years have also brought crucial insights into the functions of the AIS: how ion channels at its surface generate and shape the action potential." [12] We provide the physics and mathematics of how AIS shapes the action potential (or more precisely, we show what an important role it plays in forming AP).
Inventing AIS changed the viewpoint of neuroscience [12], a half century later, after setting up the commonly accepted model by HH. "The axon initial segment (AIS) is located at the proximal axon and is the site of action potential initiation. This reflects the high density of ion channels found at the AIS. … The summation of synaptic inputs gives rise to action potentials at the axon initial segment (AIS), a 20–60 $\mu m$ long domain located at the proximal axon/soma interface that has a high density of voltage-gated ion channels." As discussed in [13], see also their Figure 1, the structure of the Axon Initial Segment is known to the most minor details. As the illuminating investigations in 2008 [10] revealed, the AIS has very dense ion channels. That is, from an electrical point of view, those parallelized channels can be abstracted as a discrete conductance (or resistance) between the membrane and the axon. The membrane itself can be abstracted as a distributed condenser with no resistance. This model is in contrast with the viewpoint of biophysics, that the membrane plus AIS is considered a distributed element, where the capacitor and condenser cannot be separated. The current model is the result of projecting the physical processes in the resting state to the transient state. Notice the important point: "Neurons are also anatomically polarized, as they can be subdivided into a somatodendritic input domain and an axonal output domain" [9]; providing direct evidence that (unlike in HH’s model) the input and output currents (and voltage time derivatives) are independent.
Table 2: RC circuit types
The RC Integrator The RC differentiator
$V_{out}^{Integrator}=\frac{1}{RC}\int_{0}^{t}V_{in}{dt}$ $V_{out}^{Differentiator}=RC\frac{dV_{in}}{dt}$
Low Pass Filter High Pass Filter
Resting state Transient state
[Uncaptioned image] [Uncaptioned image]
(Mostly) resting state (Mostly) transient state
Ion channels in the membrane Ion channels in the AIS
In our model, the AIS gets part of the membrane, and this leads to crucial changes. (BTW, the name is slightly misleading: the AIS is part of the neuronal oscillator, and it forwards a traveling potential wave to the axon instead of belonging to it.) ’Although by definition a neuron must have an axon to assemble an AIS, the relationship between AIS assembly and axon specification in vivo has not been determined yet’ [9]. Anyhow, a neuron can be abstracted (in the disciplinary view of electricity) as comprising a resistor-like and a capacitor-like component. However, in the resting and the transient states, they are combined in different ways.
In the resting state, we can model neurons as a parallel $RC$ circuit (see the "integrator" in Table 2), where "random" currents flow in and out. Practically, the output voltage is constant (the resting potential): the ion channels and pumps are distributed over the membrane’s surface, and due to the low speed of currents, at low local voltage gradients, the AIS is not involved. Only the synaptic inputs cause a measurable gradient on the AIS. In this native state, there is no gradient; in this sense, it is similar to the case of clamped operation, where the gradient is (apparently) suppressed by the feedback. HH did not see any structural elements on the membrane, so logically, they modeled it as a distributed resistor and capacitor, which really has resemblance with a parallel $RC$ oscillator. However, they made a wrong choice of the circuit type, and their choice (likely due to inertia) was repeated in good textbooks such as ([30] Figure 3.1 or [25] Figure 1.1), and it is a commonly accepted fallacy even today [68].
In the transient state, the high-transmission ion channels in the AIS (high conductance $R$) replace the low-transmission ion channels (low conductance $R$) in the membrane’s wall, and the rush-in current flows out entirely on the AIS (for the model, in this state, we omit the low-transmission ion drain). Those two elements form a serial $RC$ circuit (see the "differentiator" in Table 2) that acts as a damped oscillator (see section 4.4) and produces a current pulse, known as AP. Neuroscience mistakenly claims it is a parallel $RC$ circuit; that statement is valid only in the resting state. However, the number and intensity of the ’resting’ ion channels in the membrane’s wall are insignificant for producing AP, so we use the approximation that only the ’transient ion channels’ in the AIS are active when generating an AP.
@@ 4.2 The oscillator model
In electric view, the process can be modeled as follows: a sudden voltage gradient (see Figure 12 middle inset) appears on the membrane (as a discrete capacitor), and a current flows out (see Figure 12 bottom inset) through the serially connected ion channel array (as a discrete resistor). The neuron forms a simple serially (not parallelly, as assumed by Hodgkin and Huxley [49] and mistakenly claimed by neurophysiology) connected $RC$ oscillator. For the physical and mathematical description, see section 4.3, furthermore, [63], section 4; for its algorithmic specifics [64], for further details [16].
Figure 12: The physical processes describing the membrane’s operation. The rushed-in $Na^{+}$ ions instantly increase the charge on the membrane, then the membrane’s capacity discharges (produces an exponentially decaying voltage derivative). The ions are created at different positions on the membrane, so they need different times to reach the AIS, where the current produces a peaking voltage derivative. The resulting voltage derivative, the sum of the two derivatives (the AIS current is outward), drives the oscillator. Its integration produces the membrane potential. When the membrane potential crosses the voltage threshold value, it switches the synaptic currents off/on.
The fundamental difference between the parallel and serial $RC$ circuits is that the parallel circuit cannot produce output voltage with opposite sign. As we discussed, the capacitive current, which by definitionem changes its direction and so generates an opposite voltage on the resistor represented by the AIS, perfectly describes the so-called "hyperpolarization". It is not the effect of a $K^{+}$ current: the resting ion channels do not have enough transmission capacity. Not knowing about the AIS leads to assuming that the resting ion channels work also as transient channels (in other words, assuming a parallel $RC$ circuit instead of the serial one, furthermore, that the output current flows through the membrane instead of the axon). Some $K^{+}$ current certainly exists; for the magnitude of currents during AP, see [15].